2. Consider again the avocado example, where demand and supply functions are Qd 160 40p Qs-50+15p Suppose a severe drought hit California, and the state government decided to subsidize farmers 40 cents for each pound of avocados produced. (Unit: Q is million pounds of avocados, and p is dollars). (1) With government subsidy, write down the functions of demand and supply. (2) What is the new equilibrium price and quantity of avocados? (Rounding to two decimal places) (3) Plot the demand and supply curves in the same graph. Label your horizontal and vertical axis properly (you can graph in either (Q, p) or (p, Q) space) Indicate the end points of each curve on each axis. Indicate the change of direction for any curve when there is government subsidy (note: the change of direction depends on whether you graph in (Q, p) or (p, Q) space)

(4) What is the price that farmers receive, and the price that consumers pay for each pound of avocados? (5) What is the percentage of subsidy pass-through to consumers? What is the percentage of subsidy that farmers retain? (6) How much in total does the state government have to subsidize the farmers? (7) Graph the above demand and supply curves in Excel Hint: choose different points of price and quantity to identify the new and old demand and supply curves; you may find it useful to make a price-quantity table first. Also, note that p and Q cannot be less than 0. (8) Suppose the 40 cents subsidy is given to consumers. Repeat questions in (1)-(7). How do your answers change?
Answer #1

Answer 1) with government subsidy, the suppliers get now p+.4, instead of per unit price p, so only supply curve shifts, demand curve unchanged

So new supply curve: Q= 50+15*(p+.4)

Unchanged demand curve: Q = 160-40*p

2) so at new equilibrium:

Demand = new supply

50+15p +6 = 160-40p

55p = 104

Thus P* = 1.89$

New quantity: from demand curve : 160-40*1.89 = 84.36 units

4) price received by farmers is 1.89+.4 = $2.29

Price paid by consumers = $1.89